OpenStudy (anonymous):

Marble (calcium carbonate) reacts with hydrochloric acid solution to form calcium chloride solution, water, and carbon dioxide. What is the percent yield of carbon dioxide if 3.65 g of the gas is collected when 10.0 g of marble reacts?

sam (.sam.):

\[CaCO_3+2HCl \rightarrow CaCl_2+H_2O+CO_2\] You have to find which is limiting first, we take 10g of marbles \[10g~marbles \times \frac{1~mol~marbles}{40+12+3(16)g~marbles} \times \frac{1~mol~CO_2}{1~mol~marbles}=0.4~mol~CO_2\] Then \[0.4~mol~CO_2 ~\times \frac{12+2(16)g~CO_2}{1~mol~CO_2}=17.6g~of~CO_2\] Based on theoretical calculations, you should get 17.6g of CO2, but it turns out that you only have 3.65g of CO2 collected, so, \[\frac{3.65g}{17.6g}\times 100=20.7 ~percent\]

OpenStudy (anonymous):

how can i find the limiting?

OpenStudy (anonymous):

sam

OpenStudy (anonymous):

help

OpenStudy (anonymous):

it's conversion... - mass of marble to moles of marble - moles of marble to moles of carbon dioxide - moles of carbon dioxide to mass of carbon dioxide sometimes you might be done here but the question also asks for percent yield, so you calculate that too (actual yield / theoretical yield)

OpenStudy (anonymous):

The problem says that a certain amount of marble reacts and says nothing about the amount of hydrochloric acid, so we can assume there is enough of the acid to react with the whole amount of marble. Thus, the amount of marble limits the amount of carbon dioxide produced.

sam (.sam.):

3.65g of CO2 is the limiting reagent because its lower than the theoretical yield which is 17.6g of CO2

OpenStudy (anonymous):

...not because it is less than theoretical yield (actual yield will almost always be smaller than theoretical yield because few reactions are perfectly efficient)

OpenStudy (anonymous):

so the acutal yield is 3.65 and the T.Y is mole of CaCO3 x atomic weight of CO2?

OpenStudy (anonymous):

CaCO3 x atomic weight of CO2 but just because the moles of CaCO3 and CO2 in this equation are the same

OpenStudy (anonymous):

is the atomic weight of CO2 is 43.99?

OpenStudy (anonymous):

i got 83%. am I right?

OpenStudy (anonymous):

AW: yes, depending on the precision of the numbers you are using (about 44) %Y: Sam seems right

OpenStudy (anonymous):

83 %?

OpenStudy (anonymous):

where are you getting 83%?

OpenStudy (anonymous):

from 3.65/4.399x100

OpenStudy (anonymous):

3.65/(.4*43.99)

OpenStudy (anonymous):

where did the 4 came from?

OpenStudy (anonymous):

no that's it... the 0.4 is moles of CO2

OpenStudy (anonymous):

from the equation above where did this came from > 40+12+3(16)g marbles

OpenStudy (anonymous):

oh sorry i thought that was 2(16), now i get it

OpenStudy (anonymous):

cool

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

you're welcome

OpenStudy (anonymous):

thanks for the medal. go chem! :-)